# MIRAGE: KV Cache Optimization through Parameter Remapping for Multi-tenant LLM Serving

### v1

@@ 7.5. Effectiveness of Our Layer Selection Strategy
The layer selection strategy is important for enabling MIRAGE to effectively pipeline GPU computation with CPU–GPU parameter transfer, thereby minimizing performance overhead. The main trade-off involves selecting the number of layers ($m$, as defined in § 5.4) to transfer from CPU to GPU when $\alpha$ layers are remapped. Using $m=\alpha+1$ reduces the amount of data transfer, while $m=\alpha+2$ better addresses data dependencies by allowing full pipelining through a double-buffering mechanism. We present an empirical evaluation of three configurations: (A) using $m=\alpha+1$ consistently, (B) using $m=\alpha+2$ consistently, and (C) a dynamic scheme that selects $m=\alpha+1$ for small $\alpha$ and switches to $m=\alpha+2$ as $\alpha$ increases.
Figure 15 presents the tail latency and throughput of serving the OPT-13b model on the Alpaca dataset using the different layer selection strategies described above. Our findings show that using $m=\alpha+2$—where transferred layers share a two-layer memory region to enable double buffering—is necessary for optimal performance. Using $m=\alpha+1$ consistently (A) can result in a reduction in $12.7\%$ throughput compared to using $m=\alpha+2$ consistently (B) or switching to $m=\alpha+2$ as $\alpha$ increases (C). The reduction in the amount of data transfer when using $m=\alpha+1$ does not outweigh the overhead caused by disruptions to the GPU execution pipeline.
@@ 7.6. When and How Many Layers to Remap?
We conduct ablation studies to explore additional design parameter choices in MIRAGE. While parameter remapping needs to be initiated only when KV cache capacity is exhausted, determining whether to “halt” remapping during off-peak periods is also important (§ 7.6.1). Moreover, more aggressive remapping can help avoid recomputation and reduce tail latency, but it may increase average latency due to increased data transfer overhead (§ 7.6.2).
@@ 7.6.1. When to halt remap?
MIRAGE monitors KV cache utilization at runtime and detects when free KV cache is available, typically during non-peak periods. For example, real-world traces show that request volume during non-peak times drops to approximately 20% of that seen during peak periods (Stojkovic et al., 2024b). To optimize performance in such periods, MIRAGE supports “halting” parameter remapping by reconfiguring the previously remapped memory space back for parameter usage. We name this mechanism Dynamic Reversion and evaluate it in this section.
Figure 16 shows the cumulative distribution function (CDF) comparison of TBT latency when serving the OPT-13b model on the Alpaca dataset, with and without the use of Dynamic Reversion. During peak load (e.g., arrival rate of 10.0 RPS), the Dynamic Reversion mechanism offers limited benefit, as parameter remapping is consistently needed to avoid recomputation. However, during non-peak periods (e.g., 1.0 RPS), Dynamic Reversion significantly improves performance, reducing P50 latency by 49% and consistently yielding lower TBT.
@@ 7.6.2. How many layers to remap?
A key design parameter in MIRAGE is the remapping percentage—the proportion of model parameter memory dynamically repurposed for KV cache. Increasing the remapping percentage enables the system to support higher request volumes during peak periods, which are difficult to predict offline. However, aggressive remapping introduces higher CPU–GPU data transfer overhead, which can interfere with the performance of concurrent requests.
We evaluate the impact of remapping percentage using OPT-13b with Alpaca dataset. As shown in Figure 17, aggressive remapping reduces tail latency under light workloads (e.g., 1.0 RPS), but increases average latency due to higher CPU–GPU data transfer overhead. In contrast, capping the remapping percentage offers a more balanced performance: it sustains low average latency across a range of workloads while still reducing tail latency under variable request rates. For instance, at 10 requests per second, the capped strategy reduces P99 latency by 58% and P50 latency by 44%.
@@ 8. Discussion and Related Work
MIRAGE is highly compatible with existing LLM optimizations. Moreover, evolving hardware trends further emphasize the need for a system like MIRAGE to fully exploit the growing CPU-GPU bandwidth. We also highlight the key differences between MIRAGE and other related memory offloading or swapping approaches.
@@
Compatibility with Other LLM Optimization Solutions: MIRAGE is designed to be highly compatible with state-of-the-art LLM inference optimizations. When the Dynamic Remapping Engine reaches its maximum remapping percentage, MIRAGE can further leverage CPU computation offloading-based methods, such as NEO (Jiang et al., 2024) and PowerInfer (Song et al., 2023), to ensure the quality of service for the online inference serving system. MIRAGE is also fully compatible with quantization methods (Lin et al., 2024; Zhao et al., 2024; Xia et al., 2024) and KV cache compression frameworks (Liu et al., 2023, 2024b, 2024a). Unlike KV cache swapping, MIRAGE requires no changes to CUDA kernels, making it compatible with advanced LLM kernels, such as Pod-Attention (Kamath et al., 2024) and Lserve (Yang et al., 2025). As a scheduler-independent memory management engine, MIRAGE supports both temporal and spatial sharing, making it compatible with a wide range of multi-tenant scheduling policies—including temporal (Bai et al., 2020; Gujarati et al., 2020; Xiao et al., 2020; Fu et al., 2024; Patke et al., 2024), spatial (Romero et al., 2021; Strati et al., 2024; Dhakal et al., 2020; Choi et al., 2022; Xu et al., 2022; Han et al., 2024a; Wang et al., 2024a), and hybrid approaches (Duan et al., 2024; Yu et al., 2025). Since MIRAGE activates parameter remapping only when GPU memory falls short for KV cache, it serves as a fallback mechanism for schedulers that mispredict memory demand under varied request arrival rates.
@@
Evolving LLM Trends: The growing popularity of multi-agent workflows (Talebirad and Nadiri, 2023; Chan et al., 2023; Li et al., 2024b; Han et al., 2024b; Kim et al., 2025) makes MIRAGE particularly well-suited for scenarios involving temporal GPU resource sharing. Furthermore, as the trend toward deploying multiple smaller-scale LLMs on a single host becomes more prevalent (Wang et al., 2024b), MIRAGE offers additional benefits by enabling efficient memory reuse across co-located models. For emerging models with smaller KV caches enabled by multi-query attention (Dubey et al., 2024), MIRAGE remains compatible by repurposing the parameter memory of inactive models to expand the available memory for active ones—offering greater flexibility than KV cache swapping based approaches.
@@
Evolving Hardware Trends: We believe the NVIDIA Grace Hopper Superchip is not just a special case but a clear demonstration of the growing trend toward increased CPU-GPU bandwidth in future hardware. NVLink, the underlying interconnect, is already mature and widely adopted. Its successor, GB200 (Nvidia, 2025b), preserves the 900 GB/s bidirectional bandwidth while scaling to 2 GPUs per node and up to 72 GPUs in an NVLink-connected rack-scale system. This architecture ensures that MIRAGE remains compatible with multi-GPU and distributed LLM inference optimizations (Sun et al., 2024; Fu et al., 2024; Yu et al., 2025). The growth in CPU performance, along with ongoing improvements in CPU DRAM bandwidth, further reinforces the trend toward high CPU–GPU bandwidth. Unlike traditional DDR memory, the LPDDRX memory used in GH200 offers significantly higher bandwidth with lower power consumption, enabling such high CPU-GPU transfer rates (Lee et al., 2018; Park et al., 2024; Li et al., 2018). As memory subsystems become increasingly critical in datacenter architectures (Maruf et al., 2023; Mahar et al., 2023), LPDDRX adoption is expected to grow, further supporting this trend.
@@
Related Works: Although prior work has explored using CPU memory as an extension of GPU memory, the focus has predominantly been on training scenarios, as seen in systems such as DeepSpeed (Rajbhandari et al., 2020; Yao et al., 2024; Ren et al., 2021; Rajbhandari et al., 2021) and others (Huang et al., 2020; Wang et al., 2022; Yuan et al., 2024; Lim et al., 2021). Training workloads are generally compute-bound, making them more tolerant of memory offloading overhead. In contrast, inference serving is latency-sensitive rather than throughput-oriented, making such memory offloading significantly more challenging. Consequently, only a few recent efforts have explored memory swapping in the context of LLM inference serving. Frameworks like FlexGen (Sheng et al., 2023) perform on-demand data migration from CPU to GPU, which can introduce delays as computations are forced to wait for data transfers to finish (Xu et al., 2024). Pie (Xu et al., 2024) was the first to explore the benefits of increased CPU-GPU bandwidth by swapping KV cache. However, its performance is affected by the bidirectional data transfer. MIRAGE stands apart from all prior work also as the first to support multi-tenant LLM inference serving.
@@ 9. Conclusion
This paper introduces MIRAGE, a Dynamic Remapping Engine that automatically and elastically repurposes model parameter memory to expand KV cache capacity for LLM inference serving, enabling low tail latency and high throughput during runtime. MIRAGE leverages the substantially increased CPU–GPU bandwidth in modern systems to overlap parameter transfers with GPU computation, minimizing the “remapping” overhead. MIRAGE can be deployed in any serving system and remains compatible with a wide range of popular multi-tenant GPU sharing and scheduling strategies. MIRAGE achieves 6.6%–86.7% improvement in throughput and 20.7%–99.3% reduction in tail latency compared to vLLM. Additionally, compared to KV cache swapping, parameter remapping incurs lower data transfer and synchronization overhead while offering greater flexibility in multi-tenant environments.

### v2

@@ 7.5. Effectiveness of Our Layer Selection Strategy
The layer selection strategy is important for enabling Oneiros to effectively pipeline GPU computation with CPU-GPU parameter transfer, thereby minimizing performance overhead. The main trade-off involves selecting the number of layers ($m$, as defined in § 5.4) to transfer from CPU to GPU when $\alpha$ layers are remapped. Using $m=\alpha+1$ reduces the amount of data transfer, while $m=\alpha+2$ better addresses data dependencies by allowing full pipelining through a double-buffering mechanism. We present an empirical evaluation of three configurations: (A) using $m=\alpha+1$ consistently, (B) using $m=\alpha+2$ consistently, and (C) a dynamic scheme that selects $m=\alpha+1$ for small $\alpha$ and switches to $m=\alpha+2$ as $\alpha$ increases. Figure 15 presents the tail latency and throughput of serving the OPT-13b model on the Alpaca dataset using the different layer selection strategies described above. Our findings show that using $m=\alpha+2$ – where transferred layers share a two-layer memory region to enable double buffering – is necessary for optimal performance. Using $m=\alpha+1$ consistently (A) can result in a reduction in $12.7\%$ throughput compared to using $m=\alpha+2$ consistently (B) or switching to $m=\alpha+2$ as $\alpha$ increases (C). The reduction in the amount of data transfer when using $m=\alpha+1$ does not outweigh the overhead caused by disruptions to the GPU execution pipeline.
@@ 7.6. When and How Many Layers to Remap?
We conduct ablation studies to explore additional design parameter choices in Oneiros. While parameter remapping needs to be initiated only when KV cache capacity is exhausted, determining whether to “halt” remapping during off-peak periods is also important (§ 7.6.1). Moreover, more aggressive remapping can help avoid recomputation and reduce tail latency, but it may increase average latency due to increased data transfer overhead (§ 7.6.2).
@@ 7.6.1. When to halt remap?
Oneiros monitors KV cache utilization at runtime and detects when free KV cache is available, typically during non-peak periods. For example, real-world traces show that request volume during non-peak times drops to approximately 20% of that seen during peak periods (Stojkovic et al., 2024b). To optimize performance in such periods, Oneiros supports “halting” parameter remapping by reconfiguring the previously remapped memory space back for parameter usage. We name this mechanism Dynamic Reversion and evaluate it in this section.
Figure 16 compares the CDF of TBT latency for OPT-13B on the Alpaca dataset, with and without Dynamic Reversion. At peak load (10.0 RPS), Dynamic Reversion offers limited benefit since remapping is consistently required to avoid recomputation. During non-peak periods (1.0 RPS), however, it improves performance substantially—reducing P50 latency by 49% and consistently lowering TBT.
@@ 7.6.2. How many layers to remap?
A key design parameter in Oneiros is the remapping percentage – the proportion of model parameter memory dynamically repurposed for KV cache. Higher percentages help sustain peak request loads that are hard to predict offline, but excessive remapping increases CPU–GPU transfer overhead, degrading concurrent request performance. We evaluate the impact of remapping percentage using OPT-13b with Alpaca dataset. As shown in Figure 17, aggressive remapping reduces tail latency under light workloads (e.g., 1.0 RPS), but increases average latency due to higher CPU-GPU data transfer overhead. Capping the remapping percentage yields a better balance—maintaining low average latency across workloads while still reducing tail latency. At 10 RPS, the capped strategy reduces P99 latency by 58% and P50 by 44%.
@@ 8. Discussion and Related Work
Compatibility with other LLM Optimization Solutions: Oneiros integrates seamlessly with state-of-the-art LLM inference optimizations. When the Dynamic Remapping Engine reaches its remapping percentage limit, it can leverage CPU offloading methods such as NEO (Jiang et al., 2025) and FlexInfer (Na et al., 2025) to maintain service quality. Unlike KV-cache swapping, Oneiros requires no CUDA kernel modifications, ensuring compatibility with advanced kernels like Pod-Attention (Kamath et al., 2024) and LServe (Yang et al., 2025). As a scheduler-independent memory engine, Oneiros supports temporal (Bai et al., 2020; Gujarati et al., 2020; Xiao et al., 2020; Fu et al., 2024; Patke et al., 2024), spatial (Romero et al., 2021; Strati et al., 2024; Dhakal et al., 2020; Choi et al., 2022; Xu et al., 2022; Han et al., 2024a; Wang et al., 2024a), and hybrid (Duan et al., 2024; Yu et al., 2025) sharing policies. By activating parameter remapping only when GPU memory is insufficient for KV cache, Oneiros serves as a fallback for schedulers that mispredict memory demand under dynamic workloads.
@@
Evolving LLM Trends: The growing popularity of multi-agent workflows (Talebirad and Nadiri, 2023; Chan et al., 2023; Li et al., 2024b; Han et al., 2024b; Kim et al., 2025) makes Oneiros well-suited for scenarios involving temporal GPU resource sharing. Furthermore, as the trend toward deploying multiple smaller-scale LLMs on a single host becomes more prevalent (Wang et al., 2024b), Oneiros offers additional benefits by enabling efficient memory reuse across co-located models.
@@
Evolving Hardware Trends: The Grace Hopper Superchip highlights a clear trend toward higher CPU-GPU bandwidth in future hardware. Its successor, GB200 (Nvidia, 2025b), preserves the 900 GB/s bidirectional bandwidth while scaling to 2 GPUs per node and up to 72 GPUs in an NVLink-connected rack-scale system. This architecture ensures that Oneiros remains compatible with multi-GPU and distributed LLM inference optimizations (Sun et al., 2024; Fu et al., 2024; Yu et al., 2025). The improvements in CPU DRAM bandwidth, further reinforces this trend. Unlike traditional DDR memory, the LPDDRX memory used in GH200 offers significantly higher bandwidth with lower power consumption, enabling such high CPU-GPU transfer rates (Lee et al., 2018; Park et al., 2024; Li et al., 2018).
Generic Applicability of Oneiros across GPU Types: Oneiros can also be deployed on earlier hardware generations, such as PCIe-based GPU systems with lower CPU-GPU bandwidth. The Remapping Controller in Oneiros dynamically determines the number of layers to remap based on the measured CPU-GPU transfer time per layer. On lower-bandwidth systems, Oneiros remains effective by adaptively remapping fewer layers (§ 5.3). In such environments, it can be further improved by leveraging CPU offloading (as discussed in § 3.1) to achieve additional performance gains.
@@
Other Related Work: Although prior work has explored using CPU memory as an extension of GPU memory, the focus has predominantly been on training scenarios, such as DeepSpeed (Rajbhandari et al., 2020; Yao et al., 2024; Ren et al., 2021; Rajbhandari et al., 2021) and others (Huang et al., 2020; Wang et al., 2022; Yuan et al., 2024; Lim et al., 2021). Inference serving is latency-sensitive rather than throughput-oriented, making memory offloading particularly challenging. FlexGen (Sheng et al., 2023) performs on-demand CPU–GPU data migration, often delaying computation as transfers complete (Xu et al., 2024). Pie (Xu et al., 2024) was the first to exploit higher CPU–GPU bandwidth via KV-cache swapping, but its performance suffers from bidirectional transfer overhead. Oneiros differs fundamentally—serving as the first system to enable efficient multi-tenant LLM inference with dynamic memory remapping.
@@ 9. Conclusion
We presented Oneiros, a Dynamic Remapping Engine that elastically repurposes model parameter memory to expand KV-cache capacity for LLM inference, achieving low tail latency and high throughput. Oneiros exploits modern CPU–GPU bandwidth to overlap parameter transfers with computation, minimizing remapping overhead. It integrates seamlessly with existing serving systems and supports diverse GPU sharing and scheduling policies.
Acknowledgement: We thank the anonymous reviewers for their valuable feedback and Dimitrios Liakopoulos for proofreading. This research was supported in part by NSF grant numbers #2326894 and #2425655, the UT ECE junior faculty start-up fund, UT iMAGiNE consortium, an award from the UT Machine Learning Lab (MLL), the AMD Chair Endowment, the Amazon Research Award, and Vista GPU Cluster through CGAI and TACC at UT Austin.